# Generalizable Radio-Frequency Radiance Fields for Spatial Spectrum Synthesis

### v2

@@ Abstract
We present GRaF, Generalizable Radio-Frequency (RF) Radiance Fields, a framework that models RF signal propagation to synthesize spatial spectra at arbitrary transmitter or receiver locations, where each spectrum measures signal power across all surrounding directions at the receiver. Unlike state-of-the-art methods that adapt vanilla Neural Radiance Fields (NeRF) to the RF domain with scene-specific training, GRaF generalizes across scenes to synthesize spectra. To enable this, we prove an interpolation theory in the RF domain: the spatial spectrum from a transmitter can be approximated using spectra from geographically proximate transmitters. Building on this theory, GRaF comprises two components: (i) a geometry-aware Transformer encoder that captures spatial correlations from neighboring transmitters to learn a scene-independent latent RF radiance field, and (ii) a neural ray tracing algorithm that estimates spectrum reception at the receiver. Experimental results demonstrate that GRaF outperforms existing methods on single-scene benchmarks and achieves state-of-the-art performance on unseen scene layouts.
@@ 1 Introduction
The evolution of wireless communication systems, such as WiFi and Sixth Generation (6G) cellular networks, has enabled transformative applications in domains such as smart cities and smart healthcare [15, 47, 26, 49]. These networks function as both communication backbones and sensing platforms. Reliable communication requires meticulous network planning to mitigate issues such as dead spots and outages [1], while sensing applications, e.g., WiFi-based localization [3, 43], depend on large volumes of high-quality data to train deep neural networks (DNNs). Traditionally, site surveys [20] are conducted to collect data for communication and sensing. However, conducting these surveys demands dense measurements at numerous locations, which is both time-consuming and labor-intensive [12, 25].
An alternative is to synthesize Radio-Frequency (RF) datasets via propagation modeling, generating spatial spectra that capture the received power from all directions around the receiver (§2.1). Given transmitter and receiver positions, RF propagation modeling estimates the received spectrum while accounting for reflection, diffraction, and scattering [24, 40]. In free space, Maxwell’s equations [35] enable accurate signal computation [55]; however, in complex real-world environments, directly solving them becomes intractable [64]. To approximate these scenarios, ray tracing simulations [61, 14, 34, 17, 19] model propagation paths using Computer-Aided Design (CAD) scene models [30, 38]. However, obtaining accurate CAD models and the high computational cost of ray tracing make such simulations both impractical and unreliable in practice.
@@ 2 Preliminary
(a) Receiver: $4\times 4$ antenna array
(b) Azimuth and elevation
Figure 2: Antenna array; each blue square is an antenna.
@@ 2.1 Spatial Spectrum
For a receiver configured as a $M\times N$ antenna array, as illustrated in Figure 2(a), the received signal is represented as a matrix $\mathbf{Y}=\big[y_{m,n}\big]_{m,n=0}^{M-1,N-1}$, where $y_{m,n}$ denotes the signal received at the $(m,n)$-th array element. Each element $y_{m,n}$ is computed as follows:
@@
where $\lambda$ is the signal wavelength, $d=\frac{\lambda}{2}$ is the element spacing, and $\alpha$ and $\beta$ are the azimuth and elevation angles of the incoming signal, respectively, as shown in Figure 2(b).
@@ 3 Related Work
3DGS [22, 60]-based spectrum synthesis methods [62, 53, 58, 63, 27] improve the training and inference times of NeRF2 [65] within a single scene. In contrast, our work focuses on achieving spatial generalization across scenes.
@@ 4 Methodology
(a) 3D spherical view
(b) 2D projection view
Figure 3: Illustration of the spatial spectrum in 3D and 2D views.
@@ 5.7 Case Study: Angle of Arrival (AoA) Estimation
The synthesized spatial spectrum can be used for downstream tasks such as Angle of Arrival (AoA) estimation for wireless localization [2]. We use an angular artificial neural network (AANN), which consists of a ResNet-50 backbone followed by an MLP head, to process the spectrum and estimate the AoA. This estimated AoA corresponds to the line-of-sight (LoS) propagation direction between the transmitter and receiver, enabling transmitter localization.
By generating synthetic training datasets, GRaF significantly reduces the data collection effort required to train the AANN. To demonstrate the benefits of the high-quality spectra synthesized by GRaF, we evaluate using the following two strategies for the training and testing stages:

### v3

@@ Abstract
We present GRaF, Generalizable Radio-Frequency (RF) Radiance Fields, a framework that models RF signal propagation to synthesize spatial spectra at arbitrary transmitter or receiver locations, where each spectrum measures signal power across all surrounding directions at the receiver. Unlike state-of-the-art methods that adapt vanilla Neural Radiance Fields (NeRF) to the RF domain with scene-specific training, GRaF generalizes across scenes to synthesize spectra. To enable this, we prove an interpolation theory in the RF domain: the spatial spectrum from a transmitter can be approximated using spectra from geographically proximate transmitters. Building on this theory, GRaF comprises two components: (i) a geometry-aware Transformer encoder that captures spatial correlations from neighboring transmitters to learn a scene-independent latent RF radiance field, and (ii) a neural ray tracing algorithm that estimates spectrum reception at the receiver. Experimental results demonstrate that GRaF outperforms existing methods on single-scene benchmarks and achieves state-of-the-art performance on unseen scene layouts. Our source code is publicly available at https://github.com/kangyang73/GRaF.
@@ 1 Introduction
The evolution of wireless communication systems, such as WiFi and Sixth Generation (6G) cellular networks, has enabled transformative applications in domains such as smart cities and smart healthcare [13, 38, 19, 40]. These networks function as both communication backbones and sensing platforms. Reliable communication requires meticulous network planning to mitigate issues such as dead spots and outages [1], while sensing applications, e.g., WiFi-based localization [3, 34], depend on large volumes of high-quality data to train deep neural networks (DNNs). Site surveys [15, 48, 49] are conducted to collect data for communication and sensing. However, conducting these surveys demands dense measurements at numerous locations, which is both time-consuming and labor-intensive [10, 18].
An alternative is to synthesize Radio-Frequency (RF) datasets via propagation modeling, generating spatial spectra that capture the received power from all directions around the receiver (§2.1). Given transmitter and receiver positions, RF propagation modeling estimates the received spectrum while accounting for reflection, diffraction, and scattering [17, 32]. In free space, Maxwell’s equations [28] enable accurate signal computation [45]; however, in complex real-world environments, directly solving them becomes intractable [57]. To approximate these scenarios, ray tracing simulations [54, 12, 27, 14] model propagation paths using Computer-Aided Design (CAD) scene models [23, 31]. However, obtaining accurate CAD models and the high computational cost of ray tracing make such simulations both impractical and unreliable in practice.
@@ 2 Preliminary
(a) 3D spherical view
(b) 2D projection view
Figure 2: Illustration of the spatial spectrum in 3D and 2D views.
@@ 2.1 Spatial Spectrum
For a receiver configured as a $M\times N$ antenna array, the received signal is represented as a matrix $\mathbf{Y}=\big[y_{m,n}\big]_{m,n=0}^{M-1,N-1}$, where $y_{m,n}$ denotes the signal received at the $(m,n)$-th array element. Each array element $y_{m,n}$ is computed as follows:
@@
where $\lambda$ is the signal wavelength, $d=\frac{\lambda}{2}$ is the element spacing, and $\alpha$ and $\beta$ are the azimuth and elevation angles of the incoming signal, respectively.
@@ 3 Related Work
3DGS [16, 53, 51]-based spectrum synthesis methods [55, 43, 50, 56, 20] improve the training and inference times of NeRF2 [58] within a single scene. In contrast, our work focuses on achieving spatial generalization across scenes.
@@ 5.7 Case Study: Angle of Arrival (AoA) Estimation
The synthesized spatial spectrum can be used for downstream tasks such as Angle of Arrival (AoA) estimation for wireless localization [2]. We use an angular artificial neural network (AANN), which consists of a ResNet-50 backbone followed by an MLP head, to process the spectrum and estimate the AoA. This estimated AoA corresponds to the line-of-sight (LoS) propagation direction between the transmitter and receiver, enabling transmitter localization. By generating synthetic training datasets, GRaF significantly reduces the data collection effort required to train the AANN. To demonstrate the benefits of the high-quality spectra synthesized by GRaF, we evaluate using the following two strategies for the training and testing stages: